Answer the questions below about the quadratic function.
f(x)= 3x^2 +24x + 51
Does the function have a minimum or maximum value?
What is the functions of the minimum or maximum value?
Where does the minimum or maximum value occur? x= ?

Answer by stanbon(75887)

Answer the questions below about the quadratic function.
f(x)= 3x^2 +24x + 51
Does the function have a minimum or maximum value?:::Minimum because
the coefficient of x^2 is positive.
-------------------------------------------
What is the function of the minimum or maximum value?:: It is the
vertex of the parabola.
-------------------------------------------
Where does the minimum or maximum value occur? x = -b/(2a) = -24/(6) = -4
f(-4) = 3(-4)^2 + 24(-4) + 51 = 48-96+51 = 3
----
Min:: (-4,3)
================
Cheers,
Stan H.
================